OpenStudy (anonymous):

Allen went to a deli to buy cheese that is sold by the pound. The scale that the deli uses is accurate within half of an ounce. According to the scale, Allen's purchase weighed exactly two pounds. Which of the following are possible actual weights for Allen's cheese? Note: There are 16 ounces in one pound. 1.87 pounds 2.10 pounds 1.96 pounds 1.97 pounds 2.04 pounds 2.03 pounds

OpenStudy (anonymous):

i was thinking that it could be 2nd 5th and 6th one

OpenStudy (tkhunny):

How did you use the "Note".

OpenStudy (anonymous):

oh i see so its the 1st 3rd 4th

OpenStudy (tkhunny):

No. Please answer my question.

OpenStudy (anonymous):

because 16 ounces is one pound so most likely its the 1st 3rd and 4th one

OpenStudy (anonymous):

thts how i used the note

OpenStudy (anonymous):

nevermind i figured it out

OpenStudy (tkhunny):

(2.00 - 1.87) lb = 0.13 lb 0.13 lb (16 oz / 1 lb) = 16*0.13 oz = 2.08 oz -- Definitely NOT the first one.